OpenStudy (anonymous):

Graph the equation and identify the y intercept 4y+3x= -12 graph the equation the y intercept is (type and ordered pair)

OpenStudy (anonymous):

y intercept is -3 ( 0 , -3 )

OpenStudy (anonymous):

4y = -12 - 3x y = -3x/4 - 3 y(0) = -3 (y intercept) (0, -3)